OpenStudy (anonymous):

How do I graph the function? State the y-intercept? y=3(2^x)

OpenStudy (anonymous):

For finding y intercept we put x = 0..

OpenStudy (anonymous):

The y-intercept occurs when x=0. It is the point (0, y). In this case, \[f(0)=3(2^0)=3\]So, your y-intercept is the point (0,3).

OpenStudy (anonymous):

Anything raised to the power 0 except 0 is 1. So: \[y = 3(2^0) = 3(1) = 3\]

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

To graph this, it is best to recognize that your function is just a variant of the basic exponential function:\[f(x)=a^x\]

OpenStudy (anonymous):

To make sense of this, first ask: how does the base "a", affect the graph? Well, the basic shape of the curve is unchanged. In fact, if f(x)=a^x, the y-intercept is always (0,1) regardless of the value "a" takes. But, the rate at which f(x) approaches zero and infinity greatly depend on the value "a" takes.

OpenStudy (anonymous):

I must add "the basic shape of the graph is unchanged" assumes a>0.